Documenta Mathematica

Band 13, 2008

V. Gritsenko, K. Hulek and G. K. Sankaran Hirzebruch-Mumford Proportionality and Locally Symmetric Varieties of Orthogonal

Type  1–19

Tom Leinster

The Euler Characteristic of a Category  21–49

Sergey Morozov Essential Spectrum of Multiparticle

Brown–Ravenhall Operators in External Field  51–79

A. Vishik, K. Zainoulline

Motivic Splitting Lemma  81–96

Jörg Winkelmann

On Tameness and Growth Conditions  97–101

V. Bach, J. Hoppe, D. Lundholm Dynamical Symmetries in

Supersymmetric Matrix  103–116

Jürgen Ritter and Alfred Weiss

Equivariant Iwasawa Theory: An Example  117–129

Guido Kings Degeneration of Polylogarithms and Special Values

of L-Functions for Totally Real Fields  131–159

Alex Kumjian, David Pask, Aidan Sims

C^{∗}-Algebras Associated to Coverings of k-Graphs  161–205

Alex Postnikov, Victor Reiner, Lauren Williams

Faces of Generalized Permutohedra  207–273

Denis Perrot Secondary Invariants for Frechet Algebras

and Quasihomomorphisms  275–363

Paul S. Muhly and Baruch Solel Schur Class Operator Functions

and Automorphisms of Hardy Algebras  365–411

Viorel Costeanu

On the 2-Typical De Rham-Witt Complex  413–452

Ciro Ciliberto, Gerard van der Geer

Andreotti–Mayer Loci and the Schottky Problem  453–504

Wiesława Nizioł

K-Theory of Log-Schemes I  505–551

Eric A. Carlen and Elliott H. Lieb Brascamp-Lieb Inequalities

for Non-Commutative Integration  553–584

Gereon Quick

Profinite Homotopy Theory  585–612

F. Déglise

Around the Gysin Triangle II.  613–675

Matthias Künzer Comparison of Spectral Sequences

Involving Bifunctors  677–737

Goro Shimura

Arithmetic of Hermitian Forms  739–774

Goro Shimura

The Critical Values of Certain Dirichlet Series  775–794

Christian Ausoni, Bjørn Ian Dundas and John Rognes Divisibility of the Dirac Magnetic Monopole

as a Two-Vector Bundle over the Three-Sphere  795–801

Timothy Logvinenko

Natural G-Constellation Families  803–823

Eva Viehmann The Global Structure of Moduli Spaces

of Polarized p-Divisible Groups  825–852